# Advanced pure tilt actuator for testing tilt-to-length coupling in space-based gravitational wave detection

### v1

@@ 1 Introduction
Based on our research on experimental testing setups for TTL coupling, we noticed that the pyramid produces a lateral displacement in the incoming light during rotation, thus affecting the final readout of the interference signal and thus influencing the calibration of the TTL coupling. To address this issue, we designed a self-aligning four-beam interferometer [17]. The feature of this device is that the conical prism does not cause lateral displacement of the incoming light when tilted. The device is constructed using a UV adhesive [18], providing compactness and significantly enhanced displacement measurement sensitivity. This makes it potentially applicable for high-precision TTL coupling experiments in the future.
@@ 3.1 Measurement principle
Once these conditions are satisfied, we can theoretically decouple the three degrees of freedom at the measurement light reflection point. As shown in Figure 4, assume three points on the same plane in space have coordinates $\mathrm{P}_{1}\left(\mathrm{x}_{1},\mathrm{y}_{1},\mathrm{z}_{1}\right),\mathrm{P}_{2}\left(\mathrm{x}_{2},\mathrm{y}_{2},\mathrm{z}_{2}\right)\text{ and }\mathrm{P}_{3}\left(\mathrm{x}_{3},\mathrm{y}_{3},\mathrm{z}_{4}\right)$, where the Z-axis is the sensitive direction. Given the plane’s rotation angle is very small (of the order of hundred microradians), we assume the coordinates remain unchanged, treating $\mathrm{x}_{1,2,3}$ and $\mathrm{y}_{1,2,3}$ as constants during the tilt, while the longitudinal displacements $\mathrm{z}_{1,2,3}$ change over time. The normal vector of the plane containing the mirror can be expressed through the cross-product of vectors between the three points:
@@ 3.3 Error analysis
APTA will be used for the calibration of TTL coupling, so it is necessary to analyze its own errors to avoid introducing them into the calibration results. The first concern is the dihedral angle error [22]. Due to the machining process, it is impossible to ensure that the three faces of the corner cube prism are perfectly perpendicular to each other, which affects the property of the constant optical path within the corner cube, thereby introducing calibration errors into the APTA. In this context, we need to account for the dihedral angle errors in the internal optical path calculations of the corner cube. Let’s consider the dihedral angle errors as $\delta_{12}$ and $\delta_{13}$, where $\delta_{12}$ is the dihedral angle error between the first and second reflective surfaces, and $\delta_{13}$ is the dihedral angle error between the first and third reflective surfaces.
@@
Figure 7: (a)The geometric structure of the pyramid and its position. Among them, $\mathrm{M}_{1}$, $\mathrm{M}_{2}$ and $\mathrm{M}_{3}$ are the three faces of the corner reflector, and their dihedral angle errors are $\delta_{12}$ and $\delta_{13}$. During the calculation, the vertex of the corner reflector is placed at the origin of the Cartesian coordinate system. (b)Additional optical path errors caused by the incidence angle when light is incident on the APTA plane at a nonperpendicular angle. $i$ is the angle of incidence of the measuring light relative to the plane of the pyramid.
@@
$\mathrm{OPD}_{\text{error }}\approx\underbrace{\left[x_{0}\delta_{12}+\sqrt{2x_{0}\delta_{13}}+\sqrt{2\delta_{12}\delta_{13}}+\left(\frac{3}{2}+\frac{x_{0}}{\sqrt{2}}\right)\delta_{12}^{2}+\delta_{13}^{2}\right]}_{\text{static term }}+\underbrace{\left[x_{0}\delta_{12}-\sqrt{2x_{0}\delta_{13}}\right]\eta}_{\text{second-order term }}+$ (8)
$\underbrace{\left[\left(1-\frac{x_{0}}{\sqrt{2}}\right)\delta_{12}^{2}-2\delta_{13}^{2}\right]\eta}_{\text{hird-order term }}+\underbrace{\left[-\frac{x_{0}}{2}\delta_{12}-\frac{x_{0}}{\sqrt{2}}\delta_{13}\right]\eta^{2}}_{\text{hird-order term }}+$
$\underbrace{\left[\left(\sqrt{2}-\frac{x_{0}}{2}\right)\delta_{12}^{2}-\left(2+\sqrt{2}x_{0}\right)\delta_{12}\delta_{13}+x_{0}\delta_{13}^{2}\right]\phi}_{\text{hird-order term }}+\underbrace{\left[-\sqrt{2}x_{0}\delta_{12}\right]\eta\phi}_{\text{hird-order term }}.$
The first term in the equation (static term) represents the additional optical path introduced when the corner cube reflector is static, due to the exit point position and its dihedral angle error. The second term (second-order term) and third term (third-order term) represent the additional optical path caused by the coupling of dihedral angle error, exit point position, and pitch deflection when the corner cube reflector undergoes pitch deflection. The fourth term (third-order term) represents the additional optical path caused by the coupling of dihedral angle error, exit point position, and horizontal deflection when the pymarid undergoes pitch deflection. The fifth term (third-order term) represents the additional optical path caused by the coupling of dihedral angle error, exit point position, horizontal deflection, and pitch deflection when the corner cube reflector undergoes pitch deflection.

### v2

@@ 1 Introduction
Based on our research on experimental testing setups for TTL coupling, we noticed that the pyramid produces a lateral displacement in the incoming light during rotation, thus affecting the final readout of the interference signal and thus influencing the calibration of the TTL coupling. To address this issue, we designed a self-aligning four-beam interferometer [17]. The feature of this device is that the conical prism does not cause lateral displacement of the incoming light when tilted. The device is constructed using a UV adhesive [18], providing compactness and significantly improved displacement measurement sensitivity. This makes it potentially applicable for high-precision TTL coupling experiments in the future.
@@ 3.1 Measurement principle
Once these conditions are satisfied, we can theoretically decouple the three degrees of freedom at the measurement light reflection point. As shown in Figure 4, assume three points on the same plane in space have coordinates $\mathrm{P}_{1}\left(\mathrm{x}_{1},\mathrm{y}_{1},\mathrm{z}_{1}\right),\mathrm{P}_{2}\left(\mathrm{x}_{2},\mathrm{y}_{2},\mathrm{z}_{2}\right)\text{ and }\mathrm{P}_{3}\left(\mathrm{x}_{3},\mathrm{y}_{3},\mathrm{z}_{4}\right)$, where the Z-axis is the sensitive direction. Given that the plane’s rotation angle is very small (of the order of a hundred microradians), we assume the coordinates remain unchanged, treating $\mathrm{x}_{1,2,3}$ and $\mathrm{y}_{1,2,3}$ as constants during the tilt, while the longitudinal displacements $\mathrm{z}_{1,2,3}$ change over time. The normal vector of the plane containing the mirror can be expressed through the cross-product of vectors between the three points:
@@ 3.3 Error analysis
APTA will be used for TTL coupling calibration, so it is necessary to analyze its own errors to avoid introducing them into the calibration results. The first concern is the dihedral angle error [22]. Due to the machining process, it is impossible to ensure that the three faces of the corner cube prism are perfectly perpendicular to each other, which affects the property of the constant optical path within the corner cube, thereby introducing calibration errors into the APTA. In this context, we need to account for the dihedral angle errors in the internal optical path calculations of the corner cube. Let’s consider the dihedral angle errors as $\delta_{12}$ and $\delta_{13}$, where $\delta_{12}$ is the dihedral angle error between the first and second reflective surfaces, and $\delta_{13}$ is the dihedral angle error between the first and third reflective surfaces.
@@
Figure 7: (a)The geometric structure of the pyramid and its position. Among them, $\mathrm{M}_{1}$, $\mathrm{M}_{2}$ and $\mathrm{M}_{3}$ are the three faces of the corner reflector, and their dihedral angle errors are $\delta_{12}$ and $\delta_{13}$. During the calculation, the vertex of the corner reflector is placed at the origin of the Cartesian coordinate system. (b)Additional optical path errors caused by the incidence angle when light is incident on the APTA plane at a non-perpendicular angle. $i$ is the angle of incidence of the measuring light relative to the plane of the pyramid.
@@
$\text{OPD}_{\text{error}}\approx$ $\underbrace{\left(x_{0}\delta_{12}+\sqrt{2}x_{0}\delta_{13}+\sqrt{2}\delta_{12}\delta_{13}+\left(\tfrac{3}{2}+\tfrac{x_{0}}{\sqrt{2}}\right)\delta_{12}^{2}+\delta_{13}^{2}\right)}_{\text{static term}}$ (8)
$+\underbrace{\left[x_{0}\delta_{12}-\sqrt{2}x_{0}\delta_{13}\right]\eta}_{\text{first-order term}}+\underbrace{\left\{\left(1-\frac{x_{0}}{\sqrt{2}}\right)\delta_{12}^{2}-2\delta_{13}^{2}\right\}\eta}_{\text{first-order term}}$
$+\underbrace{\left[\left(\sqrt{2}-\frac{x_{0}}{2}\right)\delta_{12}^{2}-\left(2+\sqrt{2}x_{0}\right)\delta_{12}\delta_{13}+x_{0}\delta_{13}^{2}\right]\phi}_{\text{first-order term}}$
$+\underbrace{\left[-\frac{x_{0}}{2}\delta_{12}-\frac{x_{0}}{\sqrt{2}}\delta_{13}\right]\eta^{2}}_{\text{second-order term}}+\underbrace{\left[-\sqrt{2}x_{0}\delta_{12}\right]\eta\phi}_{\text{second-order term}}$
The first term in the equation (static term) represents the additional optical path introduced when the corner cube reflector is static, accounting for the exit point position and its dihedral angle error. The second and third terms represent the linear coupling of the pitch angle $\eta$. However, the third term involves second-order polynomials of the dihedral angle error, indicating that if the dihedral angle deviation is small, its contribution to the optical path length change is negligible. The fourth term denotes the change in optical path length corresponding to the second-order polynomial of the pitch angle $\eta$. Similarly to the third term, its impact is negligible for small rotation angles and dihedral angle errors. The fifth term shows the linear coupling of the yaw angle $\phi$. Given that the second-order polynomial of the dihedral angle deviation is very small, the influence of the fifth term on the optical path length change is not significant. Similarly, the contribution from the sixth term may be negligible.